# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

### v2

@@ IV-C Sequential Convex Optimization Framework
We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints. The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 1) and then solve Problem 4 and Problem 3 iteratively with increasing weights on the slack variables until they reach negligible values. This framework is described in Algorithm 1.
@@
2:Initialize mean trajectory $\tau$ by solving Problem 1;
@@
4:     Solve Problem 4 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
5:     Compute new mean trajectory $\tau$ by solving Problem 3 using revised covariance trajectory $\Xi$ from previous step ;
@@ V Numerical Simulations
The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$. We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$. We set the cost weight matrices $Q_{k}=0$ and $R_{k}=I$ $\forall k\in\mathbb{Z}_{0}^{T-1}$.
@@ VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented.

### v3

@@ IV-C Sequential Convex Optimization Framework
We now describe the sequential convex optimization framework which is used to design optimal control laws for MJLS covariance steering with chance constraints. The framework requires a mean trajectory to be initialized (which is obtained by optimizing the unconstrained mean trajectory in Problem 4) and then solve Problem 7 and Problem 6 iteratively with increasing weights on the slack variables until they reach negligible values. This framework is described in Algorithm 1.
@@
2:Initialize mean trajectory $\tau$ by solving Problem 4;
@@
4:     Solve Problem 7 with last stored mean trajectory $\tau$ to get new covariance trajectory $\Xi$;
5:     Compute new mean trajectory $\tau$ by solving Problem 6 using revised covariance trajectory $\Xi$ from previous step ;
@@ V Numerical Simulations
The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$. We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$.
@@ VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented. Future work will analyze convergence behavior of the framework and extensions to semi-markov jump systems.
@@ Appendix A MJLS identities
Appendix B Discussion on independence of $x_{k}$ and $r_{k}$
In this section, we provide numerical results showing that the relationship between random variables $x_{k}$ and $r_{k}$ is neglible and they can therefore, be considered independent for all practical purposes.
We first compare $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$. A close agreement between the two would indicate our initial assumption. We consider a scalar uncontrolled ($B_{k}(i)=0$) MJLS system without noise or bias ($G_{k}(i)=c_{k}=0$) to simplify the analysis. The system parameters are given below:
$A_{k}(1)=0.8$ $\quad A_{k}(2)=1.2$ $\mathbb{P}=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $\quad\rho_{k}(i)=[0.7;0.3]$
Figure 3 shows three plots which compare the sample mean $\mu_{k}$ multiplied by $\rho_{k}(i)$ to $q_{k}(i)$ which is calculated analytically using eq. 4. The plots show high agreement between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$ indicating that the dependence between $\mathds{1}_{r_{k}=i}$ and $x_{k}$ is neglible.
Figure 3: Figure shows comparison between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$, where $\mu_{k}$ is calculated using samples and $q_{k}(i)$ is calculated analytically for both modes. The right most plot shows the comparison of sampled and analytical mean $\mu_{k}$, demonstrating that the number of samples considered are sufficient for statistical convergence.
To further demonstrate this, and also provide numerical evidence to prove our mean and covariance propagation equations are correct, we consider the following time-varying system:
$A_{1}(k)$ $=\begin{bmatrix}-0.5&1\\ -1&\sin(0.2\pi k)\end{bmatrix}$ $A_{2}(k)$ $=\begin{bmatrix}\cos(0.2\pi k)&0.5\\ 0.5&0.5\end{bmatrix}$ (36)
$B_{1}(k)$ $=\begin{bmatrix}1\\ 0\end{bmatrix}$ $B_{2}(k)$ $=\begin{bmatrix}0.5\\ 1\end{bmatrix}$
$G_{1}(k)$ $=\begin{bmatrix}1\\ 1.2\end{bmatrix}$ $G_{2}(k)$ $=\begin{bmatrix}0.2\\ 0.3\sin(0.2\pi k)\end{bmatrix}$
$T$ $=10$ $\mathbb{P}$ $=\begin{bmatrix}0.6&0.4\\ 0.5&0.5\end{bmatrix}$
We use a randomized open loop control law sampled uniformly from $[-10,10]$. The noise is $w_{k}\sim\mathcal{N}(1,1000)$ and the initial state distribution is given by $x_{0}\sim\mathcal{N}(\begin{bmatrix}20,100\end{bmatrix}^{\top},10I_{2})$. Figure 4 plots the comparison between sampled mean and covariance values with theoretical calculations using the mean and covariance propagation equations provided in this work for different number of samples. Our simulations demonstrate the numerical accuracy of the propagation equations, while also validating our assumption about $x_{k}$ and $r_{k}$.
Figure 4: Comparison of sampled mean and covariance values with theoretical values calculated using propagation equations for 10,100 and 1e5 samples.